1.You run a hospital with 100 rooms. Fixed daily cost is $901.00 which includes staff salary, property charges, maintenance etc. Variable cost per room is $19.00 which includes cleaning, equipment rentals, utility cost etc. which is incurred only when the room is full. You charge $50.00 per room per day. You sold 33.00 rooms today, how much profit/loss did you earn for today.

2 .For a table manufacturing company, selling price for a table is $21.00 per Unit, Variable cost is $10.00 per Unit, labor charge is $1.00 per Unit, rent is $782.00 per month and transportation is $10 per 10 tables. If 478.00 tables are sold in a month how much revenue company earns for that month?

3.A restaurant prepares 200.00 pizza slices and sells them at a rate of $15.00/slice. Expenses for the restaurant include raw material for pizza at $6.00 per slice, $108.00 as monthly rental and $42.00 monthly as insurance. Restaurant is open only for 25 days in a month. Today there was a party at nearby office so the demand for pizza went up to 222.00 slices. Due to lack of availability vendor could not fulfill the demand and analyzed a lost sale as $5.00 per item. How much profit could the restaurant earn today?

Please, show processes to solve them
Answer #1

1)Profit=sales-variable costs-fixed costs
=(rooms occupied*charge per room)-(rooms occupied*cost per room)-fixed cost
=(33*50)-(33*19)-901
=122
2)Sales=tables sold*selling price per table=478*21=10038

3)Daily fixed costs=total fixed costs/days
=(108+42)/25=6
Profit=revenue-variable -fixed costs
=(200*15)-(200*6)-6-(22*5)
=1684